A particle moves along the x-axis so that at time t ≥ 0 its position is given by x(t) = -t^3 + 6t^2 + 63t. Determine all intervals when the acceleration of the particle is negative.


1
Expert's answer
2021-02-10T01:57:47-0500

A particle moves along the x-axis so that at time t ≥ 0 its position is given by x(t) = -t^3 + 6t^2 + 63t. Determine all intervals when the acceleration of the particle is negative.


Solution :

"x(t) =-t^3 + 6t^2 +63t, \quad t\geq0 \\
\text{Velocity} \\
v(t) = \frac{dx}{dt} = -3t^2+12t+63 \\
\text{Acceleration} \\
a = \frac{dv}{dt} = -6t+12 \\
a<0 \quad when \quad -6t+12>0 \rightarrow t>2 \\
\text{Answer : acceleration is negative in } (2,+\infty)"